### v2

# biblio.bib

Abstract
We propose a novel framework for Network Stochastic Differential Equations (N-SDE), where each node in a network is governed by an SDE influenced by interactions with its neighbors. The evolution of each node is driven by the interplay of three key components: the node’s intrinsic dynamics (momentum effect), feedback from neighboring nodes (network effect), and a stochastic volatility term modeled by Brownian motion. Our primary objective is to estimate the parameters of the N-SDE system from high-frequency discrete-time observations. The motivation behind this model lies in its ability to analyze very high-dimensional time series by leveraging the inherent sparsity of the underlying network graph. We consider two distinct scenarios: i) known network structure: the graph is fully specified, and we establish conditions under which the parameters can be identified, considering the linear growth of the parameter space with the number of edges. ii) unknown network structure: the graph must be inferred from the data. For this, we develop an iterative procedure using adaptive Lasso, tailored to a specific subclass of N-SDE models. In this work, we assume the network graph is oriented, paving the way for novel applications of SDEs in causal inference, enabling the study of cause-effect relationships in dynamic systems. Through extensive simulation studies, we demonstrate the performance of our estimators across various graph topologies in high-dimensional settings. We also showcase the framework’s applicability to real-world datasets, highlighting its potential for advancing the analysis of complex networked systems.
Keywords: Directed graph, adaptive lasso estimation, graph scaling, topology estimation, quasi-likelihood.
1 Introduction
The study of temporal data on networks has gained considerable attention in recent years. In such models, the relationships between temporal variables are represented by a graph structure, allowing for the analysis of high-dimensional and interconnected systems. One notable example is the Network Autoregressive (NAR) model introduced in [nar], which leverages the network structure to handle ultra-high-dimensional time series. The NAR model is defined as:
$Y_{it}=\theta_{0}+\theta_{1}\sum_{j=1}^{d}\bar{a}_{ij}Y_{j(t-1)}+\theta_{2}Y_{i(t-1)}+\epsilon_{it},\quad i=1,\ldots,d,$
where $\epsilon_{it}$ is a Gaussian noise. We denote by $\bar{A}$ the normalized adjacency matrix, i.e., $\bar{A}=\bar{A}(\bar{a}_{ij})=\text{diag}(N_{1}^{-1},\ldots,N_{d}^{-1})A$ and $A=A(a_{ij})$ the true adjacency matrix with elements $a_{ij}=1$ for $i=j$ or if there is a connection between nodes $i$ and $j$, and 0 otherwise. The quantities $N_{i}$ represent the number of neighbors of node $i$. Each component of $Y$ is a node on the network. The parameters $\theta_{1}$ and $\theta_{2}$ are termed respectively the momentum (or node-effect) and network (effect) parameter. The model can be rewritten as
$\mathbf{Y}_{t}=\mathcal{T}_{0}+\mathbf{Q}\mathbf{Y}_{t-1}+\mathcal{E}_{t},$
with $\mathcal{T}_{0}=(\theta_{0},\ldots,\theta_{0})^{\prime}$, $\mathcal{E}_{t}=(\epsilon_{1t},\ldots,\epsilon_{dt})^{\prime}$, and $\mathbf{Q}=\mathbf{Q}(\theta_{1},\theta_{2})$ defined as:
$\mathbf{Q}=\theta_{1}\bar{A}+\theta_{2}I_{d\times d}.$ (1)
In these models, the graph structure is assumed to be known but the dimension $d$ is allowed to grow at a rate which is compatible with the number of observations.
The NAR model has been further extended in [gnar] to account for higher-order neighbor interactions. These discrete-time models assume a known graph structure and allow the system’s dimension $d$ to grow in relation to the sample size.
In continuous time, [ou-graph1] introduced the $d$-dimensional Graph Ornstein-Uhlenbeck (GrOU) process, a $d$-dimensional system driven by Lévy noise, defined as:
$\,\mathrm{d}Y_{t}=-\mathbf{Q}Y_{t-}\,\mathrm{d}_{t}+\,\mathrm{d}L_{t}$
where $\mathbf{Q}_{\theta}$ is a $d\times d$ matrix with values in the positive cone $S^{++}$ defined as in equation (1), and $L_{t}$ is a $d$-dimensional Lévy process. The parameter $\theta=(\theta_{1},\theta_{2})\in R^{2}$ and conditions like $\theta_{2}>0$ such that $\theta_{2}>|\theta_{1}|$, are required in order to guarantee ergodicity. The parameter have the same interpretation: $\theta_{1}$ represents the momentum effect and $\theta_{2}$ the network effect and are estimated via continuous time observations of $Y$. The same authors considered the inference under high frequency discrete-time observations in [ou-graph2]. For both models they assume two cases: the matrix $A$ is fully known and specified, or unknown, in which case a LASSO approach is used to reconstruct it. In the GrOU framework though the dimension $d$ is not allow to grow as in the NAR case.
[koike2020-glasso] considered a $d$-dimensional semimartingale $Y=(Y_{t})_{t\in[0,1]}$ with invertible variance-covariance matrix $\Sigma_{Y}=[Y,Y]_{1}$. The focus is to estimate the precision matrix $\Theta_{Y}=\Sigma_{y}^{-1}$ under the asymptotic scheme $d\to\infty$. In this context the parametric structure of $Y$ is not relevant as the focus is in the elements of $\Theta_{Y}$, moreover the sparsity of the precision matrix is addressed through a weighted graphical lasso approach.
Motivated by these developments, we introduce a new framework for Network Stochastic Differential Equations (N-SDE). This framework generalizes the discrete-time NAR model and continuous-time GrOU processes by incorporating: i) non-linear interactions between nodes through network effects and ii) stochastic volatility, which propagates dynamically across the network.
$\,\mathrm{d}X_{t}^{i}$ $=\bigg{(}\underbrace{b_{ii}(X^{i}_{t},\beta)}_{\mathsf{momentum\,\,effect}}+\sum_{j\in N_{i}}\underbrace{b_{ij}(X^{i}_{t},X^{j}_{t};\beta)}_{\mathsf{network\,\,effect}}\bigg{)}\,\mathrm{d}t+\underbrace{\sigma_{i}(X^{i}_{t},\alpha)}_{\mathsf{node\,\,volatility}}\,\mathrm{d}W^{i}_{t},$ (2)
$i=1,2,\ldots,d$. In this model, each node on the network is represented by a stochastic differential equation. The evolution of node $i$ can be affected by its previous values as well as by nonlinear interactions with its neighbors $N_{i}$. By expanding on the decomposition proposed in [nar], the terms $b_{ii}$ represent a momentum effect whereas $\sum_{j\in N_{i}}b_{ij}$ measures network effect. In our model, we further allow for a random volatility term $\sigma_{ii}$ and for a network volatility term $\sum_{j\in N_{i}}\sigma_{ij}$ which determines the volatility propagation across the network.
Our work generalizes the work of [NIPS2010] for the following linear system of SDEs:
$\,\mathrm{d}X_{t}=\sum_{j\in N_{i}}a_{ij}X_{j}^{j}\,\mathrm{d}t+\,\mathrm{d}W^{i}_{t}$
for continuous time observations and repeated samples. In their framework the graph has a given structure represented by the elements $a_{ij}$ of the adjacent matrix and the objective is to estimate the minimal time horizon needed to fully estimate the network. Similarly [bishnoi2023graph] introduced the Brownian Graph Neural Network model defined by the following SDE:
$\,\mathrm{d}X_{t}^{i}=\left(X_{t}^{i}+\frac{F_{i}}{\gamma}\right)\,\mathrm{d}t+\sqrt{\frac{2\kappa_{B}T}{\gamma}}\,\mathrm{d}W_{t}^{i}$
where $F_{i}$ is a function of both the ‘incoming to’ and ‘outgoing from’ edges for each node $X_{t}^{i}$, i.e., $F_{i}=\sum_{j:in}F_{ij}-\sum_{j:out}F_{ij}$, $T$ is a fixed time horizon, and the rest are known parameters. The goal of this approach is to represent $F_{ij}$ as a graph neural network and estimate it using deep learning methods. In a similar spirit [bergna2023graph] introduced the Graph Neural SDE model that follows:
$\,\mathrm{d}X_{t}=f_{\phi}\left(X_{t},t,\mathcal{G}\right)\,\mathrm{d}t+\sigma(X_{t},t)\,\mathrm{d}W_{t}$
where $f_{\phi}$ is a neural network with weights $\phi$ and $\mathcal{G}$ represents the graph structure. The weights $\phi$ represent the quantity of interest.
[…]
The remainder of the paper is organized as follows. Section 2 introduces the notation, the model, and the assumptions of network scaling that ensure the viability of the procedure. Section 4 considers the case of unknown network structure and derives asymptotic properties of the estimators conditionally on the estimated graph. The graph structure is estimated through an adaptive Lasso approach. Section 5 presents simulation studies to show the performance of the estimators under different graph structures and sample sizes. Finally, Section 6 presents applications to real data.
2 Network SDEs
Given a filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\geq 0},P)$ and an adapted $d$-dimensional Brownian motion $W=(W^{1},\ldots,W^{d})$, let $(X_{t})_{t\geq 0}$ be the solution of the following system of stochastic differential equations:
$\,\mathrm{d}X_{t}^{i}$ $=\bigg{(}b_{ii}(X^{i}_{t},\beta)-\sum_{j\in N_{i}}b_{ij}(X^{i}_{t},X^{j}_{t};\beta)\bigg{)}\,\mathrm{d}t+\sigma_{i}(X^{i}_{t},\alpha)\,\mathrm{d}W^{i}_{t}$ (3)
$i=1,2,\ldots d$. We further introduce the following notation do describe the graph: $N_{i}$ denotes the neighbours of node $i$ in a graph $G=(V,E)$, where $V=[d]$ is a known set of vertices, and $E$ is a fixed (i.e. non random) and known list of edges [^3] In section 4 we will consider $E$ determinstic but unknown.. We write $G_{d}=(V_{d},E_{d})$ to highlight the dependence on the dimension $d$.
The terms $b_{ij}$ and $\sigma_{ij}$ denote the drift and diffusion functions of the model. They are known functions of unknown parameters $\theta=(\alpha,\beta)$. We allow the dimension of the parameter space to grow with the size of the graph, i.e. $\pi_{d}^{\alpha}=\pi_{\alpha}(G_{d})$, $\pi_{d}^{\beta}=\pi_{\beta}(G_{d})$. We denote the total number of parameters as $\pi_{\theta}(G_{d})=\pi_{d}^{\theta}$. The parameter space is denoted with $\Theta_{d}=\Theta_{d}^{\alpha}\times\Theta_{d}^{\beta}$, a compact subset of $\mathbb{R}^{\pi_{d}^{\alpha}+\pi_{d}^{\beta}}$. We denote with $\theta_{0}\in\text{Int}\Theta_{d}$ the true value of the parameter.
We denote by $\bm{b}=(\bm{b}_{ij})$ the matrix whose elements are defined as:
$\bm{b}_{ij}=\begin{cases}b_{ii}&i=j,\,\,i=1,\ldots,d\\ b_{ij}&i=1,\ldots,d,\,\,j\in N_{i}\\ 0&\text{ otherwise}\end{cases}$
and set $\bm{\sigma}=\mathrm{diag}(\sigma_{i},i=1,\ldots,d)$. Model (3) can be rewritten in compact matrix form as
$\,\mathrm{d}X_{t}=(L_{A}\odot\bm{b})\mathbf{1}\,\mathrm{d}t+\bm{\sigma}\,\mathrm{d}W_{t}.$
where $L_{A}=I-A$, $A$ is the adjacency matrix, $\mathbf{1}=(1,\ldots,1)\in\mathbb{R}^{d}$, $\odot$ is the Hadamard (element-wise) multiplication. Denote with $\Sigma=\sigma\sigma^{\top}$.
2.1 Estimation under known graph structure
We consider discrete time observations from model (3) under usual high-frequency asymptotics, i.e., the sample path of $X$ is observed at $n+1$ equidistant discrete times $t_{i}^{n}$, such that $t_{i}^{n}-t_{i-1}^{n}=\Delta_{n}<\infty$ for $i=1,\ldots,n$ with $t_{0}^{n}=0$. We denote the discrete observations of the sample path of $X$ by $\mathbf{X}_{n}:=(X_{t_{i}^{n}})_{0\leq t_{i}\leq n}$, under the following asymptotic scheme: $\Delta_{n}\longrightarrow 0$ as $n\to\infty$, $n\Delta_{n}\to 0$, in such a way that $n\Delta_{n}\geq n^{\epsilon_{0}}$, for some $\epsilon_{0}>0$, $n\Delta^{2}_{n}\to 0$.
[…]
  1. ($\mathbf{A1}$)
    the functions $b$ and $\sigma$ are smooth (see assumption D1 (i), (iii) in [yoshida2011polynomial]), uniformly Lipschitz with polynomial growth;
  2. ($\mathbf{A2}$)
    There exists $\tau>0$ such that $\tau^{-1}\leq\Lambda_{\min}(\Sigma(x,\alpha))$, uniformly in $x$ and $\alpha$.
  3. ($\mathbf{A3}$)
[…]
  4. ($\mathbf{A4}$)
    $b(x,\beta)=b(x,\beta_{0})\text{ for }\mu_{\theta_{0}}\,a.s.\text{ all }\,x\Rightarrow\alpha=\alpha_{0}$.
    $\Sigma(x,\alpha)=\Sigma(x,\alpha_{0})\text{ for }\mu_{\theta_{0}}\,a.s.\text{ all }\,x\Rightarrow\beta=\beta_{0}$
Remark 1.
(A2) implies (D1)-(ii) of [yoshida2011polynomial]. (A4) is essentially Kessler’s identifiability condition A6. See also [KITAGAWA2014]. It implies that the random fields
$\mathbb{Y}(\alpha;\theta_{0})=-\frac{1}{2}\int_{\mathbb{R}^{d}}\left\{\mathrm{Tr}\left({\Sigma}(x,\alpha)^{-1}{\Sigma}(x,\alpha_{0})-{I}_{d}\right)+\log\frac{|{\Sigma}(x,\alpha)|}{|{\Sigma}(x,\alpha_{0})|}\right\}\mu(dx).$
$\mathbb{Y}(\beta;\theta_{0})=-\frac{1}{2}\int_{\mathbb{R}^{d}}\langle\Sigma(x,\alpha_{0})^{-1},(b(x,\beta)-b(x,\beta_{0}))^{\otimes 2}\rangle\mu(dx).$
are such that $\mathbb{Y}\neq 0$ for $\theta\neq\theta_{0}$. This, and the fact that the model is defined on a compact set imply [D3] and [D4] in [yoshida2011polynomial]. On this point see the remark in [yoshida2011polynomial, p. 462], and [uchida2012adaptive, p. 2894].
A sufficient condition for ergodicity can be obtained as in Remark 1 (i) in [uchida2012adaptive].
Further assumptions on the structure of the graph are needed:
  1. ($\mathbf{G1}$)
    Network parametrization scaling: For any $d$
    $\frac{\pi_{d}}{|E_{d}|}\leq K$
  2. ($\mathbf{G2}$)
    Graph scaling: For any $d$, and for any $\epsilon$ there exists $n_{0}$ such that for any $n>n_{0}$
    $\frac{|E_{d}|}{n\Delta_{n}}\leq\epsilon.$
Remark 2.
Conditions $\mathbf{G1}$ and $\mathbf{G2}$ control the growth of the total number of parameters as the network dimension grows in terms of the number of observations $n$. It amounts to say that each function is allowed to have an approximately constant number of parameters and the number of edges should be of the same order of the number of parameters of the model.
2.2 Properties of the QMLE estimator
We begin by analyzing the properties of the quasi-likelihood estimator under the assumption that the graph G is known and $d$ fixed, but potentially very large.
We assume we observe data generated by model (3), where the neighborhoods $N_{i}$ are known, and we consider the quasi likelihood function for $(\alpha,\beta)$
$\ell_{n}(\alpha,\beta)$ $=$ (4)
$\sum_{i=1}^{n}\left\{\frac{1}{2\Delta_{n}}\langle C^{-1}_{i-1}(\alpha),(\Delta X_{t_{i}}-b_{A,i-1}(\beta))^{\otimes 2}\rangle+\log\det C_{i-1}(\alpha)\right\}$
where $\Delta X_{t_{i}}=X_{t_{i}}-X_{t_{i-1}}$, $C_{i}(\alpha)=(\bm{\sigma}\bm{\sigma}^{\top})(X_{t_{i}};\alpha)$, and $b_{A,i-1}=(L_{A}\odot\bm{b}(X_{t_{i-1}},\beta))\mathbf{1}$. The quasi-likelihood estimator
$\hat{\theta}_{n,d}=(\hat{\alpha}_{n,d},\hat{\beta}_{n,d})\in\arg\min_{\alpha,\beta}\ell_{n}(\alpha,\beta).$ (5)
We write $\hat{\theta}_{n,d}$ so to stress the dependence of the estimator on both the sample size $n$ and the dimension of the network $d$. We may omit subscripts for ease of read. Throughout this section $\hat{\alpha}$ denotes the estimator (5).
Denote with
$\Gamma_{n}=\begin{pmatrix}\frac{1}{\sqrt{n}}\mathbf{I}_{\pi^{\alpha}}&0\\ 0&\frac{1}{\sqrt{n\Delta_{n}}}\mathbf{I}_{\pi^{\beta}}\end{pmatrix}\qquad$
the block matrix of the estimator rates and its graph-size scaled version.
In order to state our forthcoming result about a non-asymptotic error bound for estimation on a graph, we introduce regularity conditions on the contrast. Denote by $\partial_{\theta}\ell_{n}$ and $\partial^{2}_{\theta,\theta}\ell_{n}$ the gradient and Hessian matrix of $\ell_{n}$, respectively, and by $\partial_{\theta}\overline{\ell_{n}}=\Gamma_{n}\partial_{\theta}\ell_{n}$, $\partial^{2}_{\theta,\theta}\overline{\ell_{n}}=\Gamma_{n}\partial^{2}_{\theta,\theta}\ell_{n}\Gamma_{n}$ their scaled version.
  1. C$(r)$
    Regular contrast: The functions $\ell_{n}$, $\partial_{\theta}\ell_{n}$, $\partial_{\theta}^{2}\ell_{n}$ can be extended continuously to the boundary of $\Theta$ and there exist square integrable random variables $\xi_{n}$ with $\mathbb{E}\xi_{n}^{2}\leq J$, and $\mu>0$ s.t.
    $(i)\quad\max_{i\in[\pi_{d}]}\sup_{\theta:|\theta_{0}-\theta|\leq r}|\partial_{\theta_{i}}\overline{\ell_{n}}|\leq\xi_{n},\qquad$
    $(ii)\quad\inf_{\theta:|\hat{\theta}_{n}-\theta|\leq r}v^{\top}\partial^{2}\overline{\ell_{n}}(\theta)v>\mu|v|^{2}\qquad\forall v\in\mathbb{R}^{\pi_{d}},$
    for all $n$, $P_{\theta_{0}}\,a.s.$.
Assumption $C-(ii)$ is a fairly standard eigenvalue condition on the Hessian of the negative quasi-likelihood. Condition $C-(i)$ relates to the regularity of the drift and diffusion functions $\bm{b}$ and $\bm{\sigma}$. The bounding variables $\xi_{n}$ can be characterized in terms of the polynomial growth condition of such terms.
The next theorem shows how the $\ell_{2}$-error of the estimator can be controlled with high probability by quantities related to the regularity of the model, the edge parametrization and the graph scaling.
Theorem 1.
Suppose that Assumptions $\mathbf{A}$, $\mathbf{G}$, $\mathbf{C}(r/n\Delta_{n})$, hold true for some $r>0$. Then for every $\epsilon>0,d>0$, there is $n_{0}$ such that for $n>n_{0}$ we have
$|\hat{\theta}_{n}-\theta_{0}|^{2}\leq\frac{4\xi^{2}_{n}}{\mu^{2}}K\epsilon.$ (6)
with probability at least $1-C_{L}/r^{L}$, for some $L>0$.
Remark 3.
We stress the fact that the proof of the theorem relies on deriving an error bound on the estimator depending on the number of parameters. In general, given an estimator $\hat{\theta}_{n}$ the theorem could be proved under the following modified assumption:
  1. ($\mathbf{C1}^{\prime}$)
    Estimator scaling:
    $\sup_{n}\mathbb{E}|\Gamma_{n}^{-1}(\hat{\theta}-\theta_{0})|^{2}\lesssim\pi_{d}.$
Remark 4.
In [yoshida2021simplified] general simplified conditions are given for an estimator to satisfy, for any $d$,
$\sup_{n}\mathbb{E}|\Gamma_{n}^{-1}(\hat{\theta}-\theta_{0})|^{p}<\infty.$
Then, by the moment convergence in Th. 3.5, one has that
$\mathbb{E}|\Gamma_{n}^{-1}(\hat{\theta}-\theta_{0})|^{2}\to\mathbb{E}|\Delta|^{2},$
where $\Delta\sim\mathcal{N}(0,\mathcal{I}(\theta_{0})^{-1}$, and then $\mathbb{E}|\Delta|^{2}=\rm{tr}I(\theta_{0})^{-1}\leq\pi_{d}|I(\theta_{0})^{-1}|$.
3 Linear N-SDE estimator
We turn our attention to the simple but important case of a N-SDE model with a linear drift. In this model the network effect is given by a linear combinations of the parents of the node. We still allow for a non-linear diagonal diffusion term and directed edges. We remark that, for readability, we present the results for the linear case; however, they can be readily extended to linear combinations of univariate basis functions.
Suppose the drift functions take the form
$b_{ii}(X^{i},\beta)=\beta_{0i}-\beta_{ii}X^{i}\qquad b_{ij}(X^{i},X^{j};\beta)=\sum_{j\in N_{i}}\beta_{ij}X^{j}$ (7)
and the that the diffusion matrix is diagonal $\bm{\sigma}=\mathsf{diag}(\sigma_{j}(\alpha),j\in[d])$.
Following [uchida2012adaptive], it is possible to use the following adaptive estimation procedure
$\hat{\alpha}_{n}\in\arg\min_{\alpha}\mathcal{U}_{n}(\alpha)\qquad\hat{\beta}_{n}\in\arg\min_{\beta}\mathcal{V}_{n}(\hat{\alpha}_{n},\beta)$
where
$\mathcal{U}_{n}(\alpha)$ $=\frac{1}{\Delta_{n}}\sum_{i=1}^{n}\langle C^{-1}_{i-1}(\alpha),\Delta X_{t_{i}}^{\otimes 2}\rangle+\log\det C_{i-1}(\alpha)$ (8)
$\mathcal{V}_{n}(\alpha,\beta)$ $=\frac{1}{\Delta_{n}}\sum_{i=1}^{n}\langle C^{-1}_{i-1}(\alpha),(\Delta X_{t_{i}}-\Delta_{n}b_{A,i-i}(\beta)^{\otimes 2}\rangle$ (9)
We focus on the explicit form of $\hat{\beta}_{n}$ under the linearity assumption. In the case of diagonal noise, (9) can be rewritten as
$\mathcal{V}_{n}(\hat{\alpha}_{n},\beta)=\frac{1}{2\Delta_{n}}\sum_{i=1}^{n}\sum_{j=1}^{d}\frac{1}{\sigma^{2}_{j,t_{i-1}}(\hat{\alpha}_{n})}\left[\Delta X_{t_{i}}^{j}-\Delta_{n}\left(\beta_{0j}-\sum_{k\in N_{j}\cup\{j\}}\beta_{jk}X^{k}_{t_{i-1}}\right)\right]^{2}$
where $\bar{N}_{j}=N_{j}\cup\{j\}$. Let $\hat{\sigma}=\sigma(\hat{\alpha}_{n})$. The score can be computed as
$\partial_{\beta_{jl}}\mathcal{V}_{n}=-\sum_{i=1}^{n}\frac{X^{l}_{t_{i-1}}}{\hat{\sigma}^{2}_{j,t_{i-1}}}\left[\Delta X_{t_{i}}^{j}-\Delta_{n}\left(\beta_{0j}-\sum_{k\in N_{j}\cup\{j\}}\beta_{jk}X^{k}_{t_{i-1}}\right)\right].$
for $j\in[d],l\in[\bar{N}_{j}]$ (excluding the intercepts). In the case where the model has no intercepts, i.e $\beta_{0j}=0\,\forall j$, the estimating equations take the form
$\sum_{k\in\bar{N}_{j}}\beta_{jk}\sum_{i=1}^{n}\frac{X_{t_{i-i}}^{k}X_{t_{i-i}}^{l}}{\hat{\sigma}^{2}_{j,t_{i-1}}}=\frac{1}{\Delta_{n}}\sum_{i=1}^{n}\frac{\Delta X_{t_{i}}^{j}X_{t_{i-i}}^{l}}{\hat{\sigma}^{2}_{j,t_{i-1}}},\quad j\in[d],l\in[\bar{N}_{j}].$ (10)
From a statistical point of view, each neighborhood behaves as a small $|\bar{N}_{j}|$-dimensional VAR model and the estimates of the parameters in a neighborhood only depend on the neighbours (but the estimators are not independent). In particular, denote with $\beta^{\bar{N}_{j}}=(\beta_{j},\,j\in[\bar{N}_{j}])$ the sub vector of parameters related to neighborhood $N_{j}$, with $\mathbf{X}_{n}=(X_{t_{i}}^{j},i\in 0,\ldots,n-1,\,j\in[d])$ the data matrix and let $\Delta\mathbf{X}_{n}=(X_{t_{i}}^{j}-X_{t_{i-1}}^{j},i\in 1,\ldots,n,\,j\in[d])$. Similarly let $\mathbf{X}_{n}^{\bar{N}_{j}}$ the columns of $\mathbf{X}_{n}$ corresponding to $\bar{N}_{j}$. Let $\hat{\sigma}_{j,n}=(\hat{\sigma}_{t_{i}}(\hat{\alpha}_{n}),i=0,\ldots,n-1)$. We write $(\mathbf{X}^{\bar{N}_{j}})^{\otimes 2}_{n}=(X_{t_{i}}^{\otimes 2},i=0,\ldots,n-1)$. Let $\langle Y\rangle$ be the matrix defined by $\langle Y\rangle_{ij}=n^{-1}\sum_{k=1}^{n}Y_{ij,k}\,i\in[d_{1}],j\in[d_{2}]$ for $Y\in\mathbb{R}^{d_{1}\times d_{2}\times n}$ (possibly a vector).
With this notation, (10) can be rewritten as
$\left\langle\frac{(\mathbf{X}^{\bar{N}_{j}})^{\otimes 2}_{n}}{\hat{\sigma}_{j,n}^{2}}\right\rangle\beta^{\bar{N}_{j}}=\frac{1}{\Delta_{n}}\left\langle\frac{\Delta\mathbf{X}_{n}^{j}\mathbf{X}_{n}^{\bar{N}_{j}}}{\hat{\sigma}_{j,n}^{2}}\right\rangle$ (11)
where the above division is meant in a vectorized sense. The estimator $\hat{\beta}^{\bar{N}_{j}}$ can then be computed as
$\hat{\beta}^{\bar{N}_{j}}=\frac{1}{\Delta_{n}}\left\langle\frac{(\mathbf{X}_{n}^{\bar{N}_{j}})^{\otimes 2}}{\hat{\sigma}_{j,n}^{2}}\right\rangle^{-1}\left\langle\frac{\Delta\mathbf{X}_{n}^{j}\mathbf{X}_{n}^{\bar{N}_{j}}}{\hat{\sigma}_{j,n}^{2}}\right\rangle.$ (12)
The result above could be generalized to a linear drift model with diagonal diffusion term of the form
$\sigma_{j}(x^{[\bar{N}_{j}]},\alpha)=\sqrt{\alpha_{j}+(x^{j})^{2}+f(x^{[N_{j}]})},$
i.e. with a neighborhood-dependent volatility term, with $f$ bounded and non-negative.
4 Adaptive Lasso estimation of the graph structure
We now consider the case where the adjacency matrix $A$ is not known. The goal is to recover the graph structure from the data using a regularization technique.
To this aim we need to slightly modify the setup as follows. We introduce auxiliary parameters $w$ that play the role of edge weights.
We augment the parameter space as $(\theta,w)=(\alpha,\beta,w)$ with $w=\mathrm{vec}(w_{ij},1\leq i,j\leq d,\,j\neq i)$ and we recast model (3) in the following form
$\,\mathrm{d}X_{t}^{i}$ $=\bigg{(}b_{ii}(X^{i}_{t},\beta)-\sum_{j=1,j\neq i}^{d}w_{ij}b_{ij}^{\prime}(X^{i}_{t},X^{j}_{t};\beta)\bigg{)}\,\mathrm{d}t+\sigma_{i}(X^{i}_{t},\alpha)\,\mathrm{d}W^{i}_{t}$ (13)
so that the adjacency matrix $A$ of the graph is modeled as $A_{ij}=\mathds{1}(w_{ij}\neq 0)$. In this formulation the edge pattern can be recovered by applying a LASSO-type regularization to the weights $w$.
We denote by $b_{ij}(x,y;\beta,w)=w_{ij}b_{ij}^{\prime}(x,y;\beta)$, for $i\neq j$ in the same spirit of model (3). The parameter $w$ varies in the compact domain $\Theta_{w}\in\mathbb{R}^{d(d-1)}$, and we identify the entries of the vector $w$ with the extra-diagonal elements of a matrix. The true value $w_{0}\in\rm{Int}\Theta_{w}$ is such that $w_{0,ij}\neq 0$ if $A_{ij}=1$. Denote with $(\theta_{0},w_{0})$ the true parameter value.
In order to ensure model identifiability we introduce the following extension to condition $\mathbf{A1}$:
  1. ($\mathbf{A1}^{\prime}$)
    $b_{ij}(x,y;\beta)$ and $\sigma_{i}$ satisfy assumption $\mathbf{A1}$ for all $i,j$ and, for any $\beta\in\Theta_{\beta}$
    $b_{ij}(\cdot,\cdot;\beta,w)=0\,\,\forall x,y\,\,\Leftrightarrow w_{ij}=0,\quad i\neq j.$
This assumption ensures that any multiplicative constant in the model is modeled by the $w$ parameters.
We propose a two-step procedure to estimate the graph and the parameters. In the first step we obtain an initial non-regularized estimate for the both the diffusion and drift parameters based on a consistent estimator. We focus on the quasi-likelihood estimator, even though this approach can be generalized to any consistent estimator, see e.g. [de2021regularized]. We then build a penalized estimator of LSA (least squares approximation) type. Estimators of these form for diffusion-type processes have been studied in, e.g., [de2012adaptive, suzuki2020penalized, de2021regularized].
Denote with $(\tilde{\theta},\tilde{w})$ the quasi-likelihood estimator of $(\theta,w)$ given by
$(\tilde{\theta},\tilde{w})\in\arg\min_{\theta,w}\mathcal{\ell}_{n}(\theta,w)$ (14)
where $\ell_{n}$ denotes the quasi-likelihood function (4) computed with respect to the augmented model (13).
Let $\tilde{G}=\partial^{2}\ell(\tilde{\theta},\tilde{w})$ be the Hessian of $\mathcal{\ell}_{n}$ at $(\tilde{\alpha},\tilde{w})$. Let $\hat{\mathcal{G}}=\Gamma_{n}^{-1}\hat{G}$ the scaled empirical information matrix, where the scaling matrix takes into account the additional $w$ parameters and is given by
$\Gamma_{n}=\begin{pmatrix}\frac{1}{\sqrt{n}}\mathbf{I}_{\pi^{\alpha}}&0\\ 0&\frac{1}{\sqrt{n\Delta_{n}}}\mathbf{I}_{\pi^{\beta}+d^{2}}.\end{pmatrix}$
[…]
Theorem 2.
Under A’, $\Gamma_{n}(\tilde{\theta}_{n}-\theta_{0},\tilde{w}_{n}-w_{0})\Rightarrow\mathcal{N}_{\pi_{d}+d^{2}}(0,V)$.
where $V$ is the positive definite matrix representing the Fisher information of the diffusion.
We introduce the least squares type loss function
$\mathcal{F}_{n}(\theta,w):=\langle\tilde{G},(\theta-\tilde{\theta},w-\tilde{w})^{\otimes 2}\rangle+\lambda_{n}\|(w,\theta)\|_{1,\gamma(n,d)}$
where $\|\cdot\|_{1,\gamma(n,d)}$ denotes the weighted $\ell_{1}$ norm with weight vector $\gamma(n,d)=(\gamma^{\alpha}_{n},\gamma^{\beta}_{n},\gamma^{w}_{n,d})$, i.e.,
$\|(\alpha,w)\|_{1,\gamma(n,d)}=\sum_{i=1}^{\pi^{\alpha}}\gamma^{\alpha}_{n,i}|\alpha_{i}|+\sum_{i=1}^{\pi^{\beta}}\gamma^{\beta}_{n,i}|\beta_{i}|+\sum_{1\leq i,j\leq d,\,i\neq j}\gamma^{w}_{n,d,ij}|w_{ij}|$
The adaptive lasso-type estimator can be formulated as
$(\hat{\theta},\hat{w})\in\arg\min_{\theta,w}\,\mathcal{F}_{n}(\theta,w)$ (15)
Estimator (15) allows for simultaneous penalization of the graph-identifying parameters $w$ and of the non-null components in $\theta$. Denote with $s^{\alpha},s^{\beta}$ the number of non null parameters in $\alpha_{0},\beta_{0}$, respectively. For ease of notation suppose that the parameter vectors are rearranged so that the first $s^{\alpha}$ components of $\alpha_{0}$ are non-null, and similarly for $\beta$. The number of non-zero entries in $w$ corresponds to $|E|$.
Let $\bar{\gamma}_{n,d}^{w}=\max_{(i,j)\in E}\gamma^{w}_{n,d,ij}$, $\bar{\gamma}_{n}^{\alpha}=\max_{i\leq s^{\alpha}}\gamma^{\alpha}_{n,i}$, $\bar{\gamma}_{n}^{\beta}=\max_{i\leq s^{\beta}}\gamma^{\beta}_{n,i}$ the largest weights for the non-null components, and let $\check{\gamma}_{n,d}^{w}=\min_{(i,j)\notin E}\gamma^{w}_{n,d,ij}$, $\check{\gamma}_{n}^{\alpha}=\min_{i\geq s^{\alpha}}\gamma^{\alpha}_{n,i}$, $\check{\gamma}_{n}^{\beta}=\min_{i\geq s^{\beta}}\gamma^{\beta}_{n,i}$ the smallest weight for the null components.
  1. ($\mathbf{L1}$)
    Adaptive weights rates:
    $\frac{|E|\,\bar{\gamma}_{n,d}^{w}}{\sqrt{n\Delta_{n}}}=O_{p}(1)\qquad\frac{s^{\alpha}\bar{\gamma}_{n}^{\alpha}}{\sqrt{n}}=O_{p}(1)\qquad\frac{s^{\beta}\bar{\gamma}_{n}^{\beta}}{\sqrt{n\Delta_{n}}}=O_{p}(1)$
  2. ($\mathbf{L2}$)
[…]
Notice that the adaptive coefficients for $w$ may depend on both the sample size $n$ and the dimension $d$.
Remark 5.
Condition L1 depend on the sparsity of the parameter rather than on the full size of the parameter space. Under this framework we can consistently estimate N-SDEs on sparse large graphs.
A common choice for the adaptive weights is ([zou2006adaptive])
$\gamma_{n,j}^{\alpha}\propto{|\tilde{\alpha}_{n,j}|^{-\delta_{1}}},\,j\in[\pi^{\alpha}],\quad\gamma_{n,j}^{\beta}\propto{|\tilde{\beta}_{n,j}|^{-\delta_{2}}},\,j\in[\pi^{\beta}],$ (16)
$\,\gamma_{n,ij}^{w}\propto{|\tilde{w}_{n,ij}|^{-\delta_{3}}},(i,j)\in E$ (17)
where $\delta_{i}>0$. The idea is that the data-driven weights penalize more coefficients whose initial guess has small magnitude.
Theorem 3.
Under assumptions $\mathbf{A}^{\prime}$ and $\mathbf{L1}$, the Lasso estimator in (15) is consistent, i.e.
$\Gamma_{n}(\hat{\theta}_{n}-\theta_{0},\hat{w}_{n}-w_{0})=O_{p}(1).$
Let $\hat{A}$ the adjacency matrix estimator derived from (15), i.e.
$\hat{A}_{ij}=\begin{cases}\mathds{1}(\hat{w}_{ij}\neq 0)&\,i\neq j\\ 0&i=j\end{cases}\qquad 1\leq i,j\leq d.$ (18)
Let $\hat{G}=(V,\hat{E})$ the graph built by means of the estimated adjacency matrix $\hat{A}$. The next theorem shows the estimated graph is contained in the true graph with probability tending to one. In other words, no edges are falsely included.
Theorem 4.
Under assumptions $\mathbf{A}^{\prime}$ and $\mathbf{L1}$ and $\mathbf{L2}$
$P(\hat{G}\subset G)\to 1.$
Remark 6.
[koike2020-glasso] proposed a graphical lasso method for the estimation of the covariance matrix of a general semi-martingale setting. Our methodology is different because it allows the estimation of directed graph relations. As an example, see Section 5. We are able to do so because the adjacency matrix appears in the drift equations of the model.
After estimating the adjacency matrix one can build a reduced N-SDE model based on the estimated neighborhoods $\hat{N}_{i}$.
$\,\mathrm{d}X_{t}^{i}$ $=\bigg{(}b_{ii}^{\prime}(X^{i}_{t},\beta)-\sum_{j\in\hat{N}_{i}}b_{ij}^{\prime}(X^{i}_{t},X^{j}_{t};\beta)\bigg{)}\,\mathrm{d}t+\sigma_{i}(X^{i}_{t},\alpha)\,\mathrm{d}W^{i}_{t}$ (19)
and re-estimate $\theta=(\alpha,\beta)$ by quasi-likelihood and by suppressing the auxiliary parameters $w$.
5 N-SDE estimation on synthetic data
Consider the following ergodic SDE model
$\,\mathrm{d}X^{i}_{t}=\left(\mu_{i}-\sum_{j\in N_{i}\cup\{i\}}\alpha_{ij}X^{j}_{t}\right)\,\mathrm{d}t+\beta_{i}\sqrt{1+X^{2}_{t}}\,\mathrm{d}W^{i}_{t}\qquad i=1,\ldots,d.$ (20)
where $\mu_{i},\alpha_{ij}\in\mathbb{R}$ for $i\neq j$, $\alpha_{ii},\beta_{i}>0$. We test our estimation procedure on different graph configurations. In each case we focus on the capability of the model of recovering some different relevant aspect of the graph.
Erdős–Rényi graph.
In this case we are interested in estimating both the parameter values and the graph structure. Here we consider $d=10$ and the graph is represented in 1(a). We set the parameter space to be $[-10^{3},10^{3}]$ for the real valued parameters and $[0,10^{3}]$ for the non negative parameters and we used the tuning parameter $\delta=1$ for the adaptive weights and $\lambda=0.1\cdot\lambda_{\max}$. The estimates of the initial quasi likelihood estimator (14) and the lasso estimator (15) are reported in Table 2. We see in 1(b) that the adjacecny matrix estimator (18) can perfectly recover the graph.
Polymer configuration.
In this case we consider a polymer type of graph, $d=12$, following an example in [ou-graph2]. Here we introduce an important modification, i.e. the graph is oriented. All of the nodes are linked in a chain, but some nodes have double links, one per direction. The adjacency matrix is thus not symmetric. The adjacency matrix of the graph is estimated as in (18).The tuning parameter is chosen by evaluating the validation loss, and then by using the more conservative choice $\lambda_{.5se}$, which corresponds to the minimum of the validation loss plus half its standard deviation. The graph and the estimated adjacency matrix are represented in Figure 2. In this case we were able to correctly identify existing relations between nodes as well as the direction of such relations.
Stochastic block model.
In this study we aim at recovering the cluster structure of a graph. In order to test this, we consider a graph generated from a stochastic block model. Here $d=21$, the true graph is made up of three blocks of 4, 11 and 6 nodes respectively with intra-cluster connection probability $p_{in}=0.9$ and extra-cluster connection probability $p_{ex}=0.05$. The true graph and the communities are shown in 3(a). We first estimate the graph adjacency according to (18) and then use Louvain community detection algorithm to identify the clusters. The edges have been estimated by setting the penalization parameter to $\lambda_{.5se}$. The true and estimated adjacency matrices are shown in 3(b). We see that, even though the reconstructed edges do not match perfectly the true ones, our model is capable of identifying the correct cluster structure. We also show in Figure 4 the number of cluster identified as a function of the penalization parameter, compared with the validation loss.
(a) Graph representation of the neighborhood in model (20) for a Erdős–Rényi random graph
(b) Estimated and true adjacency matrices
Figure 1: Erdős–Rényi random graph
(a) Graph representation of the neighborhood in model (20) for a directed polymer graph
(b) Adjacency matrices
Figure 2: Polymer graph
(a) Graph representation of the neighborhood in model (20) for a stochastic block model graph
(b) Estimated and true adjacency matrices
Figure 3: Cluster identification in a stochastic block model
Figure 4: Clusters selected (red) and loss (blue) in a SBM as a function of the penalization parameter, on log-scale. Vertical line represents $\lambda_{0.5se}$
5.1 Empirical analysis of the error bound
We now empirically validate the results of Theorem 1. In particular, Table 1 contains a numerical computation of the mean squared error of the estimator (12), for different values of the number of edges, parameters and observation time. Empirical results show perfect agreement with the theoretical bound in (6) in terms of the expected behaviour as a function of the ratios $K$ and $\epsilon$.
$d$ $|E_{d}|$ $\pi_{d}$ $T$ $K$ $\epsilon$ Bound Mean Error
8 20 36 10 1.8 2 3.6 0.94 (0.34)
20 1 1.8 0.52 (0.16)
40 0.5 0.9 0.28 (0.08)
80 0.25 0.45 0.15 (0.04)
100 0.2 0.36 0.12 (0.04)
160 0.125 0.225 0.08 (0.02)
200 0.1 0.18 0.06 (0.02)
2000 0.01 0.018 0.007 (0.002)
16 48 80 96 1.7 0.5 1.35 0.23 (0.03)
200 0.24 0.4 0.123 (0.02)
32 100 204 200 1.45 0.5 1.2 0.23 (0.025)
Table 1: Simulation Results with different graph configurations for the estimator (12)
Par. LASSO Quasi Lik. True
$\mu_{0}$ 0.6975 1.0802 1.0
$\mu_{1}$ 0.5053 1.0445 1.0
$\mu_{2}$ 0.3421 0.7878 1.0
$\mu_{3}$ 0.3093 0.8409 1.0
$\mu_{4}$ 0.5276 1.1689 1.0
$\mu_{5}$ 0.0000 0.5910 1.0
$\mu_{6}$ 0.6798 1.2084 1.0
$\mu_{7}$ 0.6601 1.2190 1.0
$\mu_{8}$ 0.5134 1.0356 1.0
$\mu_{9}$ 0.6519 0.9590 1.0
$\alpha_{00}$ 9.2533 10.9324 10.0
$\alpha_{01}$0.0000 0.2015 0.0
$\alpha_{02}$0.0000 0.5041 0.0
$\alpha_{03}$ 3.1138 4.6660 4.5
$\alpha_{04}$$0$0.1643 0.0
$\alpha_{05}$ $0$0.2544 0.0
$\alpha_{06}$$0$0.8162 0.0
$\alpha_{07}$ 3.7750 4.7743 4.5
$\alpha_{08}$0.0000 0.2154 0.0
$\alpha_{09}$0.0000 0.2244 0.0
$\alpha_{10}$$0$0.1335 0.0
$\alpha_{11}$ 9.8537 11.2609 10.0
$\alpha_{12}$0.0000 1.0622 0.0
$\alpha_{13}$$0$0.4401 0.0
$\alpha_{14}$0.0000 0.0746 0.0
$\alpha_{15}$ 0.0000 0.2427 0.0
$\alpha_{16}$$0$0.0338 0.0
$\alpha_{17}$ 0.0000 1.4447 0.0
$\alpha_{18}$ 3.2648 4.7682 4.5
$\alpha_{19}$ 0.0000 0.3693 0.0
$\alpha_{20}$0.0000 0.3516 0.0
$\alpha_{21}$$0$0.5691 0.0
$\alpha_{22}$ 6.6206 8.8756 10.0
$\alpha_{23}$ 0.0000 0.5603 0.0
$\alpha_{24}$$0$0.7148 0.0
$\alpha_{25}$ 2.1092 3.8566 4.5
$\alpha_{26}$0.0000 0.2885 0.0
$\alpha_{27}$ 2.6212 3.4551 4.5
$\alpha_{28}$$0$0.0799 0.0
$\alpha_{29}$$0$0.3116 0.0
$\alpha_{30}$ 3.4853 4.8683 4.5
$\alpha_{31}$ 0.0000 1.0532 0.0
$\alpha_{32}$ 0.0000 0.2943 0.0
$\alpha_{33}$ 9.7901 11.1668 10.0
$\alpha_{34}$ $0$0.2725 0.0
$\alpha_{35}$$0$0.8380 0.0
$\alpha_{36}$$0$0.7394 0.0
$\alpha_{37}$0.0000 0.3025 0.0
$\alpha_{38}$ 0.0000 0.9577 0.0
$\alpha_{39}$0.0000 0.0359 0.0
$\alpha_{40}$ 0.0000 1.0301 0.0
$\alpha_{41}$$0$1.1558 0.0
$\alpha_{42}$ 0.0000 1.0767 0.0
$\alpha_{43}$$0$0.7946 0.0
$\alpha_{44}$ 8.3322 9.9380 10.0
$\alpha_{45}$0.0000 0.9905 0.0
$\alpha_{46}$ 2.9438 4.6182 4.5
$\alpha_{47}$ 4.1103 6.0158 4.5
$\alpha_{48}$$0$0.6027 0.0
$\alpha_{49}$ 0.0000 1.0776 0.0
Par. LASSO Quasi Lik. True
$\alpha_{50}$$0$0.6558 0.0
$\alpha_{51}$ 0.0000 0.4062 0.0
$\alpha_{52}$ 2.5156 3.9729 4.5
$\alpha_{53}$$0$0.7804 0.0
$\alpha_{54}$0.0000 0.1319 0.0
$\alpha_{55}$ 7.1298 9.3011 10.0
$\alpha_{56}$ 1.4615 3.3353 4.5
$\alpha_{57}$$0$0.1560 0.0
$\alpha_{58}$$0$0.2808 0.0
$\alpha_{59}$0.0000 0.2874 0.0
$\alpha_{60}$ 0.0000 0.2179 0.0
$\alpha_{61}$$0$0.2317 0.0
$\alpha_{62}$ 0.0000 0.3440 0.0
$\alpha_{63}$0.0000 0.1317 0.0
$\alpha_{64}$ 3.4087 4.5651 4.5
$\alpha_{65}$ 3.4937 4.9690 4.5
$\alpha_{66}$ 8.4420 10.2551 10.0
$\alpha_{67}$0.0000 0.1050 0.0
$\alpha_{68}$0.0000 0.0087 0.0
$\alpha_{69}$0.0000 0.3310 0.0
$\alpha_{70}$ 3.6536 4.9594 4.5
$\alpha_{71}$ 0.0000 0.2511 0.0
$\alpha_{72}$ 3.8585 5.0281 4.5
$\alpha_{73}$$0$0.1114 0.0
$\alpha_{74}$ 3.3865 4.6748 4.5
$\alpha_{75}$$0$0.1492 0.0
$\alpha_{76}$0.0000 0.2697 0.0
$\alpha_{77}$ 8.6921 10.7766 10.0
$\alpha_{78}$0.0000 0.1111 0.0
$\alpha_{79}$0.0000 0.0479 0.0
$\alpha_{80}$ 0.0000 0.2762 0.0
$\alpha_{81}$ 2.1106 3.9253 4.5
$\alpha_{82}$0.0000 0.1767 0.0
$\alpha_{83}$$0$0.6376 0.0
$\alpha_{84}$$0$0.2071 0.0
$\alpha_{85}$ 0.0000 0.5338 0.0
$\alpha_{86}$ 0.0000 0.4924 0.0
$\alpha_{87}$ 0.0000 0.2271 0.0
$\alpha_{88}$ 8.0191 9.3704 10.0
$\alpha_{89}$ 0.0000 0.3800 0.0
$\alpha_{90}$$0$0.3475 0.0
$\alpha_{91}$$0$0.5505 0.0
$\alpha_{92}$$0$1.1980 0.0
$\alpha_{93}$ 0.0000 0.9177 0.0
$\alpha_{94}$ 0.0000 0.4065 0.0
$\alpha_{95}$ 0.0000 0.3239 0.0
$\alpha_{96}$0.0000 0.7491 0.0
$\alpha_{97}$$0$1.0595 0.0
$\alpha_{98}$$0$0.6295 0.0
$\alpha_{99}$ 9.5622 10.7617 10.0
$\beta_{0}$ 1.0087 1.0101 1.0
$\beta_{1}$ 1.0159 1.0165 1.0
$\beta_{2}$1.0015 -1.0004 1.0
$\beta_{3}$ 1.0158 1.0161 1.0
$\beta_{4}$1.0098 -1.0084 1.0
$\beta_{5}$ 0.9890 0.9899 1.0
$\beta_{6}$ 1.0053 1.0057 1.0
$\beta_{7}$ 1.0191 1.0182 1.0
$\beta_{8}$ 1.0076 1.0073 1.0
$\beta_{9}$1.0097 -1.0112 1.0
Table 2: N-SDE parameter estimates in a Erdős–Rényi random graph.
6 Applications to real data
In order to test our method on real-world data, we consider high-frequency financial data. We take the component stocks of S&P100 on May 2024. Our observations closing prices during 5 minutes intervals. Accounting only for complete cases, we have $d=99$ variables $n=1596$ observations. We fit estimator (15) for a linear drift model (7) and constant diffusion. Our enlarged parameter space for $(\theta,w)$ has dimension $\pi_{d}+d^{2}=9900$, thus we are in a high-dimensional setting. The resulting graph is shown in Figure 5. Vertices are colored according to their Global Industry Classification Standard (GICS) sectors. Our graph exhibits some features that have been observed in the literature for similar types of data, see e.g. [koike2020-glasso] and [barigozzi2018power]. Our graph consists primarily of a handful of large, connected components with multiple hubs, accompanied by many small isolated components. The degree distribution is shown in . It demonstrates a heavy-tailed pattern, as the most connected nodes have a disproportionately larger number of links. The estimated networks display characteristics that are often observed in power-law graphs ([barigozzi2018power]).
Figure 5: Estimated graph for the components of S&P100 components stocks.
Figure 6: Vertex degree distribution for the S&P100 graph.
7 Conclusions.
In this paper we introduce a novel model for stochastic differential equations on networks. This model allows to deal with high-dimensional systems of SDEs, by modeling the interactions between the series by means of a graph. The novelty in this model lies in the possibility of having general non-linear relations in both the drift interactions and the volatility, as well as directed graph relations. Our contribution is two-fold. On the one hand we provide a form of non-asymptotic control on the estimation error that takes into account the graph scaling in relation of the observation time as well as the graph parametrization. Roughly speaking, this tackles the questions on how much time one needs to observe the graph and how many parameters one can have for each edge in order to have a reliable estimate. On the other hand we analyze a LASSO-based graph estimation procedure, that allows to recover the graph based on the temporal information. We validate our findings by means empirical studies on simulated and real data.
8 Proofs
Proof of Theorem 1..
We prove that the bound holds true on the event $\{|\Gamma_{n}^{-1}(\hat{\theta}_{n}-\theta_{0})|\leq r\}$. This event, under assumptions A, has probability at least $1-C_{L}/r^{L}$ due to the results in [yoshida2011polynomial]. See, e.g., [yoshida-nonlin] formula (2.14).
On the event $\{|\Gamma_{n}^{-1}(\hat{\theta}_{n}-\theta_{0})|\leq r\}$, one has that $\{|\hat{\theta}_{n}-\theta_{0}|\leq r/\sqrt{n\Delta_{n}}\}$ ando so we can apply the inequalities in Assumption $\mathbf{C}(r/n\Delta_{n})$.
First, by Taylor expansion and Cauchy-Schwartz inequality we have that
$|\ell_{n}(\hat{\theta}_{n})-\ell_{n}(\theta_{0})|$ $\leq\left|\int_{0}^{1}\partial_{\theta}\ell_{n}(\theta_{0}+u(\hat{\theta}_{n}-\theta_{0}))\cdot(\hat{\theta}_{n}-\theta_{0})\mathrm{d}u\right|$
$\leq\int_{0}^{1}|\partial_{\theta}\overline{\ell_{n}}(\theta_{0}+u(\hat{\theta}_{n}-\theta_{0}))|\mathrm{d}u\,|\Gamma_{n}^{-1}(\hat{\theta}_{n}-\theta_{0})|$
$\leq\xi_{n}\sqrt{\pi_{d}}|\Gamma_{n}^{-1}(\hat{\theta}_{n}-\theta_{0})|.$
Moreover,
$|\ell_{n}(\theta_{0})-\ell_{n}(\hat{\theta}_{n})|$ $=\left|\int_{0}^{1}(1-u)\left\langle\partial^{2}_{\theta\theta}\ell_{n}(\hat{\theta}+u(\theta_{0}-\hat{\theta}_{n})),(\hat{\theta}_{n}-\theta_{0})^{\otimes 2}\right\rangle\mathrm{d}u\right|$
$\geq\frac{\mu}{2}|\Gamma_{n}^{-1}(\hat{\theta}_{n}-\theta_{0})|^{2}.$
Putting everything together,
$\sqrt{n\Delta_{n}}|(\hat{\theta}_{n}-\theta_{0})|\leq|\Gamma_{n}^{-1}(\hat{\theta}_{n}-\theta_{0})|\leq 2\frac{\xi_{n}\sqrt{\pi_{d}}}{\mu}.$
By the assumptions G, we then have
$|\hat{\theta}_{n}-\theta_{0}|^{2}$ $\leq 4\frac{\xi_{n}^{2}{\pi_{d}}}{\mu^{2}n\Delta_{n}}\leq\frac{4\xi_{n}^{2}}{\mu^{2}}\frac{|E_{d}|}{n\Delta_{n}}\frac{\pi_{d}}{|E_{d}|}\leq\frac{4\xi_{n}^{2}}{\mu^{2}}K\epsilon.$
Proof of Theorem 3..
We follow similar steps as [de2021regularized]. In order to prove the consistency we notice that
00 0 $\geq\mathcal{F}_{n}(\hat{\theta},\hat{w})-\mathcal{F}_{n}(\theta_{0},w_{0})$
$=\langle\tilde{G},(\hat{\theta}-\theta_{0},\hat{w}-w_{0})^{\otimes 2}\rangle+2\langle\tilde{G},(\hat{\theta}-\theta_{0},\hat{w}-w_{0})\otimes(\tilde{\alpha}-\theta_{0},\tilde{w}-w_{0})\rangle$
$\quad+\lambda_{n}(\|(\hat{\theta},\hat{w})\|_{1,\gamma(n,d)}-\|(\theta_{0},w_{0})\|_{1,\gamma(n,d)})$
$\geq\|\tilde{\mathcal{G}}^{-1}\|^{-1}|\Gamma_{n}^{-1}(\hat{\theta}-\theta_{0},\hat{w}-w_{0})|^{2}$
$\quad-2\|\tilde{\mathcal{G}}\||\Gamma_{n}^{-1}(\hat{\theta}-\theta_{0},\hat{w}-w_{0})||\Gamma_{n}^{-1}(\tilde{\theta}-\theta_{0},\tilde{w}-w_{0})|$
$\quad-\lambda_{n}\left(\frac{|E|\,\bar{\gamma}_{n,d}^{w}}{\sqrt{n\Delta_{n}}}+\frac{s^{\alpha}\bar{\gamma}_{n}^{\alpha}}{\sqrt{n}}+\frac{s^{\beta}\bar{\gamma}_{n}^{\beta}}{\sqrt{n\Delta_{n}}}\right)|\Gamma_{n}^{-1}(\hat{\theta}-\theta_{0},\hat{w}-w_{0})|.$
Hence we get
$|\Gamma_{n}^{-1}(\hat{\theta}-\theta_{0},\hat{w}-w_{0})|$
$\quad\leq\|\tilde{\mathcal{G}}^{-1}\|\left[\|\tilde{\mathcal{G}}\||\Gamma_{n}^{-1}(\tilde{\theta}-\theta_{0},\tilde{w}-w_{0})|+\left(\frac{|E|\,\bar{\gamma}_{n,d}^{w}}{\sqrt{n\Delta_{n}}}+\frac{s^{\alpha}\bar{\gamma}_{n}^{\alpha}}{\sqrt{n}}+\frac{s^{\beta}\bar{\gamma}_{n}^{\beta}}{\sqrt{n\Delta_{n}}}\right)\right]$
$=O_{p}(1)$
because of Theorem 2 and assumption L.
Proof of Theorem 4.
Denote by $w^{\bullet}$ the subvector of $w$ corresponding to the null entries of the true parameter $w_{0}$. This means that $w^{\bullet}_{ij}=0\Leftrightarrow(i,j)\notin E$, and $(\hat{w}^{\bullet}=0)\subset(\hat{G}\subset G)$. Then it suffices to show that
$P(\hat{w}^{\bullet}\neq 0)\to 0.$
The proof then goes along the classical route of a selection consistency results. See e.g. [de2021regularized]. ∎
[…]